# Classifying locally distinguishable sets: No activation across bipartitions

### v1

@@ III Non-activable and activable product states
Now considering $|{\mathcal{\xi}}_{12}^{\pm}\rangle_{A}|0\rangle_{A},$ we get $\quad\left\langle 1+2\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1-2\right\rangle_{A}\langle 0|0\rangle_{B}=0,\quad$ i.e., $\left\langle 1\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1\right\rangle-\left\langle 2\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|2\right\rangle=0$. Thus, $m^{a}_{11}=m^{a}_{22}$. For the states $|{\mathcal{\xi}}_{23}^{\pm}\rangle_{A}|1\rangle_{A}$ we finally get $m^{a}_{11}=m^{a}_{22}=m^{a}_{33}.$ Therefore, ${\mathcal{M}_{A}^{m}}^{\dagger}{\mathcal{M}_{A}^{m}}$ $=\operatorname{diag}\left(\delta_{0},\gamma,\gamma,\gamma\right).$ If possible let us assume that $\delta_{0}\neq 0$ and $\gamma\neq 0$. Then after Alice’s measurement, Bob should do a nontrivial operation on his own system according to Alice’s result. We denote $\mathcal{M}_{B}^{m}$ as Bob’s operator. As we discussed above, by choosing suitable pair of states we can conclude that all the off-diagonal element of ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}$ is equal to 0. Similarly for the diagonal element as we have discussed above, if we take $|0\rangle_{A}|{\mathcal{X}}_{01}^{\pm}\rangle_{B},\;|0\rangle_{A}|{\mathcal{X}}_{23}^{\pm}\rangle_{B},\;|1\rangle_{A}|{\mathcal{X}}_{12}^{\pm}\rangle_{B}$ we finally get $m^{b}_{00}=m^{b}_{11}=m^{b}_{22}=m^{b}_{33}.$ Therefore ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}$ is propotional to the identity operator, i.e., ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}=\lambda_{0}I$, which is trivial operator and this contradicts our assumption. So, either $\delta_{0}=0$ or $\gamma=0.$ Notice that this result also suggests that these states cannot be distinguished if Bob goes first. Now it is clear that if Alice goes first with a diagonal operator i.e., $\delta_{0}=\gamma=1$, then the above set of states cannot be distinguished. So, Alice has to do non-trivial measurement first and this only happens when any one of $\delta_{0}$, $\gamma$ not equal to zero. For that Alice only has two outcome measurement operators: $\quad{\mathcal{M}_{A}^{1}}^{\dagger}\mathcal{M}_{A}^{1}=\operatorname{diag}(1,0,0,0)$ and ${\mathcal{M}_{A}^{2}}^{\dagger}\mathcal{M}_{A}^{2}$ = $I-{\mathcal{M}_{A}^{1}}^{\dagger}\mathcal{M}_{A}^{1}$ $=\operatorname{diag}(0,1,1,1)$, see Fig. 2. If ‘$1$’ clicks, Bob is able to distinguish the left states by projecting onto $\left|0\pm 1\right\rangle$ and $\left|2\pm 3\right\rangle$. If ‘$2$’ clicks, it isolates the remaining states. It is then Bob’s turn to do measurement. Following the method we used above, we can similarly prove that Bob’s measurement must be ${\mathcal{M}_{B}^{1}}^{\dagger}\mathcal{M}_{B}^{1}=\operatorname{diag}(1,0,0,0)$ and ${\mathcal{M}_{B}^{2}}^{\dagger}\mathcal{M}_{B}^{2}$ $=\operatorname{diag}(0,1,1,1)$. The process will repeat a finite number of times and for each measurement outcomes for both parties the set $\mathcal{S}_{1}$ transforms only to a distinguishable set. This implies the fact that if the set is distinguishable (local) then for all possible nontrivial measurements, it is impossible to transform the set into an indistinguishable one. In other words, the set $\mathcal{S}_{1}$ is not activable through orthogonality-preserving LOCC. Hence we complete the proof. ∎
@@
[…]
For bipartite systems the variation with respect to hidden kind of nonlocality is very limited as $k$ can have only one value ($k=1$). So, there are only two types of structures, one is activable and the other is non-activable. Both classes are weaker in the sense of nonlocality because they are not locally indistinguishable class after all.
However, multipartite Hilbert spaces provide some interesting results which cannot be seen for bipartite Hilbert spaces. More generally, for the task of activation of nonlocality the multipartite Hilbert space provides some broader view than bipartite cases. For example, in the tripartite scenario there exists a set of states, $\mathcal{S}$ which is not activable when all three parties are spatially separated, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S})=0.$, but the same set of states might be activable when two parties perform some joint operation(s), i.e., $\mathcal{H}^{\text{LOCC}}_{2}(\mathcal{S})$ may not be zero subrata2024 . Consequently, a question arises: Is it feasible to construct a tripartite set for which the activation of nonlocality by LOCC is precluded across every bipartition? Such a class would, in essence, constitute the ‘worst case’ from the perspective of non-locality activation. The subsequent findings furnish appropriate support for this aforementioned concept. Consider the set $\mathcal{S}_{2}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}\in\mathbf{\mathbf{C}^{4}\otimes\mathbf{C}^{2}\otimes\mathbf{C}^{2}}$, given by,
@@
where, $|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{A},\;|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{B},\;$
$|{\mathcal{Y}}_{ij}^{\pm}\rangle_{C}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{C}$ and $|{\mathcal{\xi}}_{k}\rangle_{A}=|k\rangle_{A},\;|{\mathcal{X}}_{k}\rangle_{B}=|k\rangle_{B}$
$|{\mathcal{Y}}_{k}\rangle_{C}=|k\rangle_{C}$.
@@
The set $\mathcal{S}_{2}$ does not possess any activable nonlocality in tripartition $A|B|C$ as well as in all bipartition under orthogonality-preserving LOCC. i.e., $(i)\;\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{2})=0$ and $(ii)\;\mathcal{H}^{\text{LOCC}}_{2}(\mathcal{S}_{2})=0.$
@@
(i) We begin by noting that given any multipartite set, if it does not contain any activable nonlocality across all bipartitions then it becomes obvious that the set also does not contain any activable nonlocality in multi-partitions. This is because in bipartitions the operations are stronger than that of the multi-partitions. For example, in our context if we consider a bipartition then, two parties can perform joint measurements but in a tripartition such a possibility is absent. Clearly, the operations in bipartitions can be much stronger than that of a tripartition. Thus, we concentrate only on proving the second part of the proposition.
(ii) We need to prove that $\mathcal{S}_{2}$ is not activable in all bipartitions. First, we consider the case A|BC. In A|BC, the states belong to a $\mathbf{\mathbf{C}^{4}\otimes\mathbf{C}^{4}}$ Hilbert space and they have the same forms as the states of the set $\mathcal{S}_{1}$. So, by Proposition 1, the set of states $\mathcal{S}_{2}$ is non-activable in A|BC bipartition.
For the bipartitions B|AC and C|AB, the states of the set $\mathcal{S}_{2}$ belong to $\mathbf{\mathbf{C}^{2}\otimes\mathbf{C}^{8}}$. Now, it is known that a set of product states in $\mathbf{\mathbf{C}^{2}\otimes\mathbf{C}^{n}}$ is always locally distinguishable BennettUPB1999 . Moreover, LOCC is not sufficient to create entanglement from product states. So, it is impossible to activate nonlocality from the states of the set $\mathcal{S}_{2}$ in B|AC and C|AB bipartitions. Hence $\mathcal{H}^{\text{LOCC}}_{2}(\mathcal{S}_{2})=0.$
@@
Let us not concentrate on the particular structure of tripartite product states, given in (2). Instead, we consider any tripartite orthogonal product states in $\mathbf{\mathbf{C}^{4}\otimes\mathbf{C}^{2}\otimes\mathbf{C}^{2}}$ such that these states mimic the similar forms like the states of (1) in $\mathbf{\mathbf{C}^{4}\otimes\mathbf{C}^{4}}$ bipartition. Then, from the aforesaid proof technique it depicts that for such a set the activable nonlocality is 0 across all bipartitions.
@@
Next, we want to discuss about a hierarchy among the multipartite locally distinguishable sets. For this purpose, we first consider the bipartite set $\mathcal{S}_{3}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}_{i=1}^{10}\in\mathbf{\mathbf{C}^{6}\otimes\mathbf{C}^{6}}$, where
[…]
@@
$\mathcal{M}_{B}^{3}:=P\left[\ket{\mathbf{0}+\mathbf{1}+\mathbf{2}+\mathbf{3}+\mathbf{4}+\mathbf{5}}_{B}\right],$
@@
$\mathcal{M}_{A}^{3}:=P[\ket{\mathbf{3}}_{A}],\mathcal{M}_{A}^{4}:=\mathbb{I}-\left(\mathcal{M}_{A}^{1}+\mathcal{M}_{A}^{2}+\mathcal{M}_{A}^{3}\right)\},$
to distinguish between these four states. This concludes the local discrimination protocol for the set $\mathcal{S}_{3}$. In the following, we will demonstrate a protocol to activate nonlocality without entanglement from this set.
@@
$\left\{\begin{array}[]{c}|\mathbf{3}\rangle_{A}|\mathbf{0}-\mathbf{1}\rangle_{B},|\mathbf{5}\rangle_{A}|\mathbf{1}-\mathbf{2}\rangle_{B},\\ |\mathbf{4-5}\rangle_{A}|\mathbf{0}\rangle_{B},|\mathbf{3-4}\rangle_{A}|\mathbf{2}\rangle_{B},\\ |\mathbf{3+4+5}\rangle_{A}|\mathbf{0}+\mathbf{1}+\mathbf{2}\rangle_{B}\end{array}\right\}$
@@
$\left\{\begin{array}[]{c}|\mathbf{0}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{2}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{1-2}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{0-1}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{0+1+2}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B},\\ |\mathbf{3}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{5}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{4-5}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{3-4}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{3+4+5}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\end{array}\right\}$
@@
$\left\{\begin{array}[]{c}|\mathbf{0}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{2}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{1-2}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{0-1}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{0+1+2}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\end{array}\right\}$
@@
$\left\{\begin{array}[]{c}|\mathbf{3}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{5}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{4-5}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{3-4}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{3+4+5}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\end{array}\right\}$
It is evident that, for each instance of Alice’s measurement, specified by the set $\{\mathcal{K}^{A}_{1},\;\mathcal{K}^{A}_{2}\}$; the five post-measurement states, contingent upon $\mathcal{K}^{B}_{1}$ clicking, yield the celebrated unextendable product basis (UPB) BennettPB1999 ; BennettUPB1999 in $\mathbf{C^{3}\otimes C^{3}}$. Also, the post-measurement states for each case of Alice’s measurement $\mathcal{K}^{A}_{1},\;\mathcal{K}^{A}_{2}$ when $\mathcal{K}^{B}_{2}$ clicks form the same UPB. See Fig. 3. It has been well established that UPB is locally indistinguishable Divincinzo ; BennettUPB1999 . So, the set $\mathcal{S}_{3}$ is activable by orthogonality-preserving LOCC. Hence, this completes the proof. ∎
[…]
Towards a hierarchy: We consider the set $\mathcal{S}_{4}=\left\{\left|\phi_{i}\right\rangle_{AB}\otimes\left|0\right\rangle_{C},\left|\phi_{i}\right\rangle_{AB}\otimes\left|1\right\rangle_{C}\right\}_{i=1}^{10}\in\mathbf{C}^{6}\otimes\mathbf{C}^{6}\otimes\mathbf{C}^{2}$, where $\left\{\left|\phi_{i}\right\rangle_{AB}\right\}_{i=1}^{10}=\mathcal{S}_{3}$. Now, consider all bipartitions of the tripartite system. For the bipartition $AB|C$, the total Hilbert space is $\mathbf{C}^{2}\otimes\mathbf{C}^{36}$, and due to the limited dimension of subsystem $C$, the set remains non-activable in this cut bandhyopadhyay201 . However, for the bipartitions $A|BC$ and $B|AC$, the set becomes activable. This follows directly from Proposition 3.
@@
Let us now highlight the contrasting structures of the sets $\mathcal{S}_{2}$ and $\mathcal{S}_{4}$. The set $\mathcal{S}_{2}$ is a tripartite ensemble of orthogonal quantum states that is initially locally distinguishable and remains non-activable in all bipartitions. In contrast, the set $\mathcal{S}_{4}$ is activable in certain bipartitions (but not in every bipartition). This structural difference reveals a clear separation in the degrees of hidden nonlocality for $\mathcal{S}_{2}$ and $\mathcal{S}_{4}$.

### v2

@@ III Non-activable and activable product states
Now considering $\mathbf{|{\mathcal{\xi}}_{12}^{\pm}\rangle_{A}|0\rangle_{A}},$ we get $\mathbf{\qquad\left\langle 1+2\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1-2\right\rangle_{A}\langle 0|0\rangle_{B}}=0,\quad$ i.e. $\mathbf{\left\langle 1\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1\right\rangle-\left\langle 2\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|2\right\rangle}=0$. Thus, $m^{a}_{11}=m^{a}_{22}$. For the states $\mathbf{|{\mathcal{\xi}}_{23}^{\pm}\rangle_{A}|1\rangle_{A}}$ we finally get $m^{a}_{11}=m^{a}_{22}=m^{a}_{33}.$ Therefore, ${\mathcal{M}_{A}^{m}}^{\dagger}{\mathcal{M}_{A}^{m}}$ $\operatorname{diag}\left(\delta_{0},\gamma,\gamma,\gamma\right).$ If possible let us assume that $\delta_{0}\neq 0$ and $\gamma\neq 0$. Then after Alice’s measurement, Bob should do a nontrivial operation on his own system according to Alice’s result. We denote $\mathcal{M}_{B}^{m}$ as Bob’s operator. As we discussed above, by choosing suitable pair of states we can conclude that all the off-diagonal element of ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}$ is equal to 0. Similarly for the diagonal element as we have discussed above, if we take $\mathbf{|0\rangle_{A}|{\mathcal{X}}_{01}^{\pm}\rangle_{B},\;|0\rangle_{A}|{\mathcal{X}}_{23}^{\pm}\rangle_{B},\;|1\rangle_{A}|{\mathcal{X}}_{12}^{\pm}\rangle_{B}}$ we finally get $m^{b}_{00}=m^{b}_{11}=m^{b}_{22}=m^{b}_{33}.$ Therefore ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}$ is propotional to the identity operator, i.e., ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}=\lambda_{0}I$, which is trivial operator and this contradicts our assumption. So, either $\delta_{0}=0$ or $\gamma=0$. Notice that this result also suggests that these states cannot be distinguished if Bob goes first. Now it is clear that if Alice goes first with a diagonal operator i.e., $\delta_{0}=\gamma=1$, then the above set of states cannot be distinguished. So, Alice has to do non-trivial measurement first and this only happens when any one of $\delta_{0}$, $\gamma$ not equal to zero. For that Alice only has two outcome measurement operators: $\quad{\mathcal{M}_{A}^{1}}^{\dagger}\mathcal{M}_{A}^{1}=\operatorname{diag}(1,0,0,0)$ and ${\mathcal{M}_{A}^{2}}^{\dagger}\mathcal{M}_{A}^{2}$ = $I-{\mathcal{M}_{A}^{1}}^{\dagger}\mathcal{M}_{A}^{1}$ $\operatorname{diag}(0,1,1,1)$, see Fig. 2. If ‘$1$’ clicks, Bob is able to distinguish the left states by projecting onto $\mathbf{\left|0\pm 1\right\rangle}$ and $\mathbf{\left|2\pm 3\right\rangle}$. If ‘$2$’ clicks, it isolates the remaining states. It is then Bob’s turn to do measurement. Following the method we used above, we can similarly prove that Bob’s measurement must be ${\mathcal{M}_{B}^{1}}^{\dagger}\mathcal{M}_{B}^{1}=\operatorname{diag}(1,0,0,0)$ and ${\mathcal{M}_{B}^{2}}^{\dagger}\mathcal{M}_{B}^{2}$ $\operatorname{diag}(0,1,1,1)$. The process will repeat a finite number of times and for each measurement outcomes for both parties the set $\mathcal{S}_{1}$ transforms only to a distinguishable set. This implies the fact that if the set is distinguishable (local) then for all possible nontrivial measurements, it is impossible to transform the set into an indistinguishable one. In other words, the set $\mathcal{S}_{1}$ is not activable through orthogonality-preserving LOCC. Hence we complete the proof. ∎
@@
Figure 2: (Color online) Representation of product states in ${\mathbb{C}}^{4}\otimes{\mathbb{C}}^{4}$. Tile indices correspond to consecutively ordered basis states of set $\mathcal{S}_{1}$, while tile colors indicate compatible measurement setups for both parties. $\mbox{M}_{i}^{j}={\mathcal{M}_{i}^{j}}^{\dagger}\mathcal{M}_{i}^{j}$; $i=\mbox{A, B}$, $j=1,2$ (this is also maintained in other figures unless explicitly stated).
For bipartite systems the variation with respect to a hidden kind of non-locality is very limited, as $k$ can have only one value ($k=1$). So, there are only two types of structures, one is activable and the other is non-activable. Both classes are weaker in the sense of non-locality because they are not locally indistinguishable class after all. Here we have provided only non-activable structure as it is going to be useful further in the multipartite scenario. However, for further research, one can also think about constructing activable scenarios in two-ququad systems.
Multipartite Hilbert spaces provide some interesting results which cannot be seen for bipartite Hilbert spaces. More generally, for the task of activation of non-locality the multipartite Hilbert space provides some broader view than bipartite cases. For example, in the tripartite scenario there exists a set of states $\mathcal{S}$, which is not activable when all three parties are spatially separated, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S})=0$, but the same set of states might be activable when two parties perform some joint operation(s), i.e., $\mathcal{H}^{\text{LOCC}}_{2}(\mathcal{S})$ may not be zero subrata2024 . Consequently, a question arises: Is it feasible to construct a tripartite set for which the activation of non-locality by LOCC is precluded across every bipartition? Such a class would, in essence, constitute the ‘worst case’ from the perspective of non-locality activation. The subsequent findings furnish appropriate support for this aforementioned concept. Consider the set $\mathcal{S}_{2}\subseteq\mathbb{C}^{4}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, given by,
@@
where, $\mathbf{|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{A},\;|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{B}},\;$
$\mathbf{|{\mathcal{Y}}_{ij}^{\pm}\rangle_{C}=\left(\frac{|i\rangle\pm|j\rangle}{\sqrt{2}}\right)_{C}}$ and $\mathbf{|{\mathcal{\xi}}_{k}\rangle_{A}=|k\rangle_{A},\;|{\mathcal{X}}_{k}\rangle_{B}=|k\rangle_{B}}$
$\mathbf{|{\mathcal{Y}}_{k}\rangle_{C}=|k\rangle_{C}}$.
@@
The set $\mathcal{S}_{2}$ does not possess any activable non-locality in tripartition $A|B|C$ as well as in all bipartition under orthogonality-preserving LOCC, i.e. $(i)\;\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{2})=0$ and $(ii)\;\mathcal{H}^{\text{LOCC}}_{2}(\mathcal{S}_{2})=0.$
@@
(i) We begin by noting that given any multipartite set, if it does not contain any activable non-locality across all bi-partitions then it becomes obvious that the set also does not contain any activable non-locality in multi-partitions. This is because in bi-partitions the operations are stronger than that of the multi-partitions. For example, in our context if we consider a bipartition then, two parties can perform joint measurements but in a tri-partition such a possibility is absent. Clearly, the operations in bi-partitions can be much stronger than that of a tri-partition. Thus, we concentrate only on proving the second part of the proposition.
(ii) To prove that $\mathcal{S}_{2}$ is not activable in all bi-partitions, consider the case A|BC. Here the states belong to a $\mathbb{C}^{4}\otimes\mathbb{C}^{4}$ and they have the same forms as the states of the set $\mathcal{S}_{1}$. So by Proposition 1, the set of states $\mathcal{S}_{2}$ is non-activable in A|BC bi-partition.
For the bipartitions B|AC and C|AB, the states of the set $\mathcal{S}_{2}$ belong to $\mathbb{C}^{2}\otimes\mathbb{C}^{8}$. Now, it is known that a set of product states in $\mathbb{C}^{2}\otimes\mathbb{C}^{n}$ is always locally distinguishable BennettUPB1999 . Moreover, LOCC is not sufficient to create entanglement from product states. So, it is impossible to activate non-locality from the states of the set $\mathcal{S}_{2}$ in B|AC and C|AB bipartitions. Hence $\mathcal{H}^{\text{LOCC}}_{2}(\mathcal{S}_{2})=0$. ∎
@@
Let us not concentrate on the particular structure of tripartite product states, given in (2). Instead, we consider any tripartite orthogonal product states in $\mathbb{C}^{4}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ such that these states mimic the similar forms like the states of (1) in $\mathbb{C}^{4}\otimes\mathbb{C}^{4}$ bipartition. Then, from the aforesaid proof technique it depicts that for such a set the activable non-locality is 0 across all bi-partitions.
@@
Next, we want to discuss about a hierarchy among the multipartite locally distinguishable sets. For this purpose, we first consider the bipartite set $\mathcal{S}_{3}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}_{i=1}^{10}\in\mathbb{C}^{6}\otimes\mathbb{C}^{6}$, where
$\begin{aligned} &\left|\phi_{1}\right\rangle_{AB}=|\mathbf{0}\rangle_{A}|\mathbf{0}-\mathbf{1}+\mathbf{4}-\mathbf{5}\rangle_{B}\\ &\left|\phi_{2}\right\rangle_{AB}=|\mathbf{2}\rangle_{A}|\mathbf{1}-\mathbf{2}+\mathbf{5}-\mathbf{3}\rangle_{B}\\ &\left|\phi_{3}\right\rangle_{AB}=|\mathbf{1-2}\rangle_{A}|\mathbf{0}-\mathbf{4}\rangle_{B}\\ &\left|\phi_{4}\right\rangle_{AB}=|\mathbf{0-1}\rangle_{A}|\mathbf{2}-\mathbf{3}\rangle_{B}\\ &\left|\phi_{5}\right\rangle_{AB}=|\mathbf{0+1+2}\rangle_{A}|\mathbf{0}+\mathbf{1}+\mathbf{2}+\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\\ &\left|\phi_{6}\right\rangle_{AB}=|\mathbf{3}\rangle_{A}|\mathbf{0}-\mathbf{1}+\mathbf{4}-\mathbf{5}\rangle_{B}\\ &\left|\phi_{7}\right\rangle_{AB}=|\mathbf{5}\rangle_{A}|\mathbf{1}-\mathbf{2}+\mathbf{5}-\mathbf{3}\rangle_{B}\\ &\left|\phi_{8}\right\rangle_{AB}=|\mathbf{4-5}\rangle_{A}|\mathbf{0}-\mathbf{4}\rangle_{B}\\ &\left|\phi_{9}\right\rangle_{AB}=|\mathbf{3-4}\rangle_{A}|\mathbf{2}-\mathbf{3}\rangle_{B}\\ &\left|\phi_{10}\right\rangle_{AB}=|\mathbf{3+4+5}\rangle_{A}|\mathbf{0}+\mathbf{1}+\mathbf{2}+\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\end{aligned}$ (3)
@@
[…]
@@
[…]
to distinguish between these four states. This concludes the local discrimination protocol for the set $\mathcal{S}_{3}$. In the following, we will demonstrate a protocol to activate non-locality without entanglement from this set.
@@
$\left\{\begin{array}[]{c}|\mathbf{3}\rangle_{A}|\mathbf{0}-\mathbf{1}\rangle_{B},|\mathbf{5}\rangle_{A}|\mathbf{1}-\mathbf{2}\rangle_{B},\\ |\mathbf{4-5}\rangle_{A}|\mathbf{0}\rangle_{B},|\mathbf{3-4}\rangle_{A}|\mathbf{2}\rangle_{B},\\ |\mathbf{3+4+5}\rangle_{A}|\mathbf{0}+\mathbf{1}+\mathbf{2}\rangle_{B}\end{array}\right\}.$
@@
$\left\{\begin{array}[]{c}|\mathbf{0}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{2}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{1-2}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{0-1}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{0+1+2}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B},\\ |\mathbf{3}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{5}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{4-5}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{3-4}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{3+4+5}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\end{array}\right\}.$
@@
$\left\{\begin{array}[]{c}|\mathbf{0}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{2}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{1-2}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{0-1}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{0+1+2}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\end{array}\right\},$
@@
$\left\{\begin{array}[]{c}|\mathbf{3}\rangle_{A}|\mathbf{4}-\mathbf{5}\rangle_{B},|\mathbf{5}\rangle_{A}|\mathbf{5}-\mathbf{3}\rangle_{B},\\ |\mathbf{4-5}\rangle_{A}|\mathbf{4}\rangle_{B},|\mathbf{3-4}\rangle_{A}|\mathbf{3}\rangle_{B},\\ |\mathbf{3+4+5}\rangle_{A}|\mathbf{3}+\mathbf{4}+\mathbf{5}\rangle_{B}\end{array}\right\}.$
It is evident that, for each instance of Alice’s measurement, specified by the set $\{\mathcal{K}^{A}_{1},\;\mathcal{K}^{A}_{2}\}$; the five post-measurement states, contingent upon $\mathcal{K}^{B}_{1}$ clicking, yields the celebrated unextendable product basis (UPB) BennettPB1999 ; BennettUPB1999 in $\mathbb{C}^{3}\otimes\mathbb{C}^{3}$. Also, the post-measurement states for each case of Alice’s measurement $\mathcal{K}^{A}_{1},\;\mathcal{K}^{A}_{2}$ when $\mathcal{K}^{B}_{2}$ clicks form the same UPB. See Fig. 3. It has been well established that UPB is locally indistinguishable Divincinzo ; BennettUPB1999 . So, the set $\mathcal{S}_{3}$ is activable by orthogonality-preserving LOCC. Hence, this completes the proof. ∎
Figure 3: (Color online) Tiling diagram for the states in $\mathbf{\mathcal{S}_{3}}$. The outlined region indicates the support of Alice’s and Bob’s measurement outcomes, resulting in post-measurement states, contained in a UPB subspace. $\mbox{M}_{i}^{j}={\mathcal{K}_{j}^{i}}$; $i=\mbox{A, B}$, $j=1,2$
Towards a hierarchy: We consider the set $\mathcal{S}_{4}=\left\{\left|\phi_{i}\right\rangle_{AB}\otimes\left|\mathbf{0}\right\rangle_{C},\left|\phi_{i}\right\rangle_{AB}\otimes\left|\mathbf{1}\right\rangle_{C}\right\}_{i=1}^{10}\in\mathbb{C}^{6}\otimes\mathbb{C}^{6}\otimes\mathbb{C}^{2}$, where $\left\{\left|\phi_{i}\right\rangle_{AB}\right\}_{i=1}^{10}=\mathcal{S}_{3}$. Now, consider all bipartitions of the tripartite system. For the bipartition $AB|C$, the total Hilbert space is $\mathbb{C}^{36}\otimes\mathbb{C}^{2}$, and due to the limited dimension of subsystem $C$, the set remains non-activable in this cut bandhyopadhyay201 . However, for the bipartitions $A|BC$ and $B|AC$, the set becomes activable. This follows directly from Proposition 3.
@@
Let us now highlight the contrasting structures of the sets $\mathcal{S}_{2}$ and $\mathcal{S}_{4}$. The set $\mathcal{S}_{2}$ is a tripartite ensemble of orthogonal quantum states that is initially locally distinguishable and remains non-activable in all bipartitions. In contrast, the set $\mathcal{S}_{4}$ is activable in certain bipartitions (but not in every bipartition). This structural difference reveals a clear separation in the degrees of hidden non-locality for $\mathcal{S}_{2}$ and $\mathcal{S}_{4}$.